Determine the values of the specified properties at each of the following conditions (a) For Refrigerant 134a at T= 60°C and v = 0.072 m3/kg, determine p in kPa, and h in kJ/kg p= h= (b) For ammonia at p = 20bar and v = 0.005 m3/kg, determine Tin kPa kJ/kg and u in kJ/kg oC kJ/kg -10°C and u T= u= (c) For Refrigerant 22 at T and v in m3/kg. 200 k /kg, determine p in bar bar m3/kg v=
Answer #1

Comment below in case of any doubt.

classmate Date Uusing Table 4-13 SupuheatdR13a at Ta 6oC ond: 0.072 m3) p 320+ (0.072- 0-81)x (2100-32o 0-66110-0.08It P 362

classmate Date Page = 0.0007606 m3/ K 0.0652 mk = 33.27 kJlk ラ1 +/200-33.27 )x(0-0652-о-o007606 1) = 0-000-7606 R33-02-33-27